# Polarity transitions induced by symmetry-breaking outer boundary heat flux in rapidly rotating dynamos

### v1

@@ 4.1 Evolution of a density disturbance in two-component convection
$P^{T}=Ra_{\ell,V}^{T}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}u_{y}\,\Theta\,\mbox{d}y\mbox{d}z,\quad P^{C}=Ra_{\ell,V}^{C}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}u_{y}\,\gamma\,\mbox{d}y\mbox{d}z.$ (74)
The integrals in (74) are computed at $x=0$ for the limits ±$\pm 20$ in $(y,z)$. (The integrals of $u_{y}\Delta T_{0}$ and $u_{y}\Delta C_{0}$ vanish). Since thermal buoyancy is not expected to contribute more than 25% of the total buoyancy power for Earth’s core (Lister & Buffett, 1995), the geodynamo can exist below the threshold for polarity reversals for $Ra_{\ell,H}/Ra^{T}_{\ell}>0.5$ (table 4), which must in turn correspond to large lower-mantle heat flux variations $q^{\star}$ of $O(10)$ as per figure 14(d). For relatively small horizontal buoyancy, the dynamo would be comfortably placed in the dipole-dominated regime without polarity transitions. In short, for nearly invariant vertical Rayleigh numbers of composition and temperature, the magnitude of the horizontal Rayleigh number would determine whether the dynamo operates in a non-reversing dipolar state, a reversing state or a multipolar state.
SL $\varLambda$ $f^{T}\%$ $Ra^{C}_{\ell,V}$ $Ra^{T}_{\ell,V}$ $Ra_{\ell,H}$ $-\omega_{A,V}^{2}/\omega_{M}^{2}$ $\omega_{A,H}^{2}/\omega_{M}^{2}$ $Ra_{\ell,H}/Ra^{T}_{\ell}$
@@
Table 4: Calculation of the relative intensity of horizontal buoyancy in two-component linear magnetoconvection for states where the slow MAC waves disappear. The fraction of thermal power $f^{T}$ is defined in (73); $Ra^{C}_{\ell,V}$, $Ra^{T}_{\ell,V}$ are the local vertical compositional and thermal Rayleigh numbers respectively. In addition, $Ra_{\ell,H}$ is the local horizontal Rayleigh number. The resultant local thermal Rayleigh number $Ra_{\ell}^{T}=Ra^{T}_{\ell,V}+Ra_{\ell,H}$. The parameters are $E_{\eta}=2\times 10^{-5}$ and $t=0.01$.
Figure 15 shows the complementarity between the vertical buoyancies of composition and temperature (both of which are taken to be constant) and the horizontal buoyancy of temperature in suppressing the slow MAC waves. In figure 15(a), the squares of the fundamental frequencies and the slow MAC wave frequency are plotted against the relative horizontal buoyancy, whereas in figure 15(b), $\omega_{A}^{2}$ is decomposed into its three parts.
Figure 15: Variation of the squares of frequencies with relative horizontal buoyancy in two-component magnetoconvection. The dotted vertical line in (a) corresponds to $|\omega_{A}|\approx|\omega_{M}|$, when the slow wave frequency $\omega_{s}$ goes to zero. Panel (b) gives the decomposition of $\omega_{A}^{2}$ into its three parts consisting of the vertical buoyancies of composition and temperature (${\omega^{C}_{A,V}}^{2}$, ${\omega^{T}_{A,V}}^{2}$) and the horizontal buoyancy of temperature ($\omega_{A,H}^{2}$). The parameters used are $E_{\eta}=2\times 10^{-5}$, $\varLambda=200$ and $f_{T}=10$% at time $t/t_{\eta}=10^{-2}$.
@@ 5 Concluding remarks
The present study investigates the dipole–multipole transition through the analysis of MHD wave motions in rapidly rotating dynamos subject to inhomogeneous heat flux at the outer boundary. The regime in focus is that of low inertia, where the nonlinear inertial force is small relative to the Coriolis force not only on the length scale of the planetary core but also on the characteristic length scale of convection. The dipole–multipole transition corresponds to the approximate parity between $|\omega_{M}|$ and $|\omega_{A}|$, where $|\omega_{M}|$ is based on the peak field intensity and $|\omega_{A}|$ is a resultant buoyancy frequency derived from the vertical and horizontal (lateral) buoyancy frequencies. A dipole-dominated strong-field dynamo of mean square field intensity $O(1)$ reaches a polarity transition by the progressive increase of horizontal buoyancy induced by an equatorially anti-symmetric heat flux heterogeneity at the outer boundary, which acts not by breaking the symmetry of the fluid columns but by suppressing the slow MAC waves. Polarity reversals lie in a narrow range of horizontal buoyancies between the dipolar and multipolar regimes. An equatorially symmetric heat flux variation does not induce the transition even at high values of the dimensionless heterogeneity $q^{\star}$ considered here since the mean temperature gradient at the equator is unaffected for a fixed vertical buoyancy. A boundary heterogeneity consisting of comparable magnitudes of symmetric and anti-symmetric variations induces the polarity transition at a value of $q^{\star}$ which is of the same order as that for the transition induced by a purely anti-symmetric variation.
@@
The present study proposes a dynamical constraint on the upper bound of the heat flux heterogeneity based on the complementarity between the vertical and horizontal buoyancies in suppressing the slow waves, noted in both the linear magnetoconvection and nonlinear dynamo models. In a single-component (thermal) buoyancy-driven dynamo, a dipole solution may be possible for any value in the range 0–0.75 of the relative horizontal buoyancy, $Ra_{\ell,H}/Ra_{\ell}$ (figure 14(c)). However, the fact that compositional buoyancy is dominant in the two-component geodynamo, together with the known order of magnitude of the peak field intensity in the inertia-free limit, suggest $Ra_{\ell,H}/Ra_{\ell}>0.5$, corresponding to $q^{\star}$ of $O(10)$ at which an axial dipole can exist. From the point of view of global mantle convection (Olson et al., 2015) and the regional stratification of the outer core (Mound et al., 2019), the maximum variation of heat flux at Earth’s CMB could be higher than the mean heat flux. As suggested by the present two-component magnetoconvection model (§4.1), a numerical upper bound for the heterogeneity in Earth’s CMB heat flux may be obtained from two-component dynamo simulations using the heterogeneity pattern derived from the seismic velocity variation in the lowermost mantle (Masters et al., 1996). A lower bound for the heterogeneity may also be obtained from dynamo simulations, based on the minimum variation needed to obtain the hemispherical (east–west) variability in the high-latitude magnetic flux in Earth (Sahoo & Sreenivasan, 2020) or the mantle-induced heat flux heterogeneity at the inner core boundary (Sreenivasan & Gubbins, 2011).
@@
For nearly invariant vertical Rayleigh numbers of composition and temperature, values of the relative horizontal buoyancy much lower than that needed to suppress the slow MAC waves in the core would place the dynamo in a non-reversing dipolar state. This provides a way to explain the existence of geomagnetic superchrons, the long periods in Earth’s history without polarity reversals. That said, even for relative horizontal buoyancies $>0.5$, the ratio of symmetric to anti-symmetric heterogeneity in heat flux at Earth’s CMB can vary in geological time. For example, Courtillot & Olson (2007) suggest a relation between superchrons and the formation of mantle plumes, which control the heat flow across the CMB. If a plume forms near the equator, the heat flux at the CMB is dominated by an equatorially symmetric heat flux pattern, promoting a dipole-dominated non-reversing state. On the other hand, if a plume forms far from the equator, the resulting anti-symmetric heat flux pattern suppresses the slow MAC waves and produces a regime conducive to reversals.
@@ Appendix A Dynamo with outer boundary heat flux derived from seismic tomography
Figure 16: Evolution of dipole colatitude with time (measured in units of the magnetic diffusion time) for the dynamo subject to heterogeneous outer boundary heat flux derived from the seismic shear wave velocity in Earth’s lower mantle at at $q^{\star}=13$. The other parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
@@
Figure 17: Section plots at height $z=0.2$ below the equator showing $\omega_{A}^{2}$ (a, c) and $|\omega_{M}^{2}/\omega_{A}^{2}|$ (b, d) for the composite outer boundary heat flux heterogeneity derived from the seismic shear wave velocity in Earth’s lowermost mantle. Two values of the heterogeneity are considered, $q^{\star}=10$ (a, b) and $q^{\star}=13$ (c, d). The other parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
@@ Appendix B Spectral coefficients for the perturbations of temperature and composition
$g_{3}$ $=\frac{\partial^{2}\hat{\Theta}}{\partial{t^{2}}}\big{|}_{t=0}=d_{2}\bigg{(}\beta_{y}^{T}\frac{k_{y}}{k_{z}}-\beta_{z}\bigg{)},$
$g_{4}$ $=\frac{\partial^{3}\hat{\Theta}}{\partial{t^{3}}}\big{|}_{t=0}=d_{3}\bigg{(}\beta_{y}^{T}\frac{k_{y}}{k_{z}}-\beta_{z}\bigg{)},$
$g_{5}$ $=\frac{\partial^{4}\hat{\Theta}}{\partial{t^{4}}}\big{|}_{t=0}=d_{4}\bigg{(}\beta_{y}^{T}\frac{k_{y}}{k_{z}}-\beta_{z}\bigg{)}.$
@@
$q_{3}$ $=\frac{\partial^{2}\hat{\gamma}}{\partial{t^{2}}}\big{|}_{t=0}=d_{2}\beta_{y}^{C}\frac{k_{y}}{k_{z}},$
$q_{4}$ $=\frac{\partial^{3}\hat{\gamma}}{\partial{t^{3}}}\big{|}_{t=0}=d_{3}\beta_{y}^{C}\frac{k_{y}}{k_{z}},$
$q_{5}$ $=\frac{\partial^{4}\hat{\gamma}}{\partial{t^{4}}}\big{|}_{t=0}=d_{4}\beta_{y}^{C}\frac{k_{y}}{k_{z}}.$

### v2

@@ 4.1 Evolution of a density disturbance in two-component magnetoconvection
$P^{T}=Ra_{\ell,V}^{T}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\bm{u}\cdot\Theta\hat{\bm{e}}_{y}\,\mbox{d}y\mbox{d}z,\quad P^{C}=Ra_{\ell,V}^{C}\int_{-\infty}^{\infty}\int_{-\infty}^{\infty}\bm{u}\cdot\gamma\hat{\bm{e}}_{y}\,\mbox{d}y\mbox{d}z.$ (76)
The integrals in (76) are computed at $x=0$ for the limits $\pm 20$ in $(y,z)$ (The integrals of $u_{y}\Delta T_{0}$ and $u_{y}\Delta C_{0}$ vanish).
4.2 Complementarity of vertical and horizontal buoyancies in two-component convection
Figure 16: Variation of the squares of frequencies with relative horizontal buoyancy in two-component magnetoconvection. The dotted vertical line in (a) corresponds to $|\omega_{A}|\approx|\omega_{M}|$, when the slow wave frequency $\omega_{s}$ goes to zero. Panel (b) gives the decomposition of $\omega_{A}^{2}$ into its three parts consisting of the vertical buoyancies of composition and temperature (${\omega^{C}_{A,V}}^{2}$, ${\omega^{T}_{A,V}}^{2}$) and the horizontal buoyancy of temperature ($\omega_{A,H}^{2}$). The parameters used are $E_{\eta}=2\times 10^{-5}$, $B^{2}_{peak}=200$ and $f^{T}=10$% at time $t/t_{\eta}=10^{-2}$.
In §2.4, the complementarity between vertical and horizontal buoyancies at the state of vanishing slow MAC waves was noted. In two-component convection, a complementarity exists between the vertical buoyancies of composition and temperature and the horizontal buoyancy of temperature in suppressing the slow MAC waves (figure 16). Figure 16(a) shows the variation of the squares of the fundamental frequencies and the slow MAC wave frequency with relative horizontal buoyancy, whereas figure 16(b) presents the decomposition of $\omega_{A}^{2}$ into its three constituent parts. While the compositional Rayleigh number $Ra^{C}_{\ell,V}$ is held constant (see below), the thermal Rayleigh number is determined by the power ratio $f^{T}$, given by (75). In figure 16(a), $f^{T}$ is set to 10% and the horizontal buoyancy is progressively increased. When the resultant $\omega_{A}$ obtained from (68) matches $\omega_{M}$, the slow MAC waves disappear.
The analysis in figure 16 is then performed for a range of $f^{T}$, and the results are summarized in table 4.
4.2.1 A constraint on the lateral heat flux variation in Earth’s lower mantle
In one-component convective dynamos, the heterogeneity factor $q^{*}$ that causes the suppression of slow MAC waves may take values of O(1)O(1)$O(10)$; see figure 15(d). In two-component convection, we aim to place an upper bound on $q^{*}$ for Earth using plausible values of the peak magnetic field intensity and the Rayleigh number in the core. For the dipole-dominated regime given by $|\omega_{C}|>|\omega_{M}|>|\omega_{A}|$, the square of the peak value of the scaled magnetic field should take values of $O(10^{2})$ (Sreenivasan and Maurya, 2021; Varma and Sreenivasan, 2022) while the mean square value of the field could be $O(1)$, as suggested by observations (Gillet et al., 2010). Furthermore, if strong compositional buoyancy in the core by itself maintains the dynamo in a dipolar state below the threshold for polarity transitions, a value of $Ra^{C}_{\ell,V}$ of $O(10^{3})$ is reasonable (Majumder et al., 2024). For Earth, this value of the local Rayleigh number corresponds to a vertical Rayleigh number $O(10^{3})$ times the critical Rayleigh number for onset of nonmagnetic convection, essential for reproducing the observed polar circulation of 0.6–0.9 yr-1 (Hulot et al., 2002) in the low-inertia geodynamo (see Majumder and Sreenivasan, 2023). Now, using the complementarity of the buoyancy frequencies (68), the value of the relative horizontal buoyancy $Ra_{\ell,H}/Ra^{T}_{\ell}$ for polarity transitions based on the condition of vanishing slow MAC waves, $|\omega_{A}|\approx|\omega_{M}|$, may be obtained for a range of thermal power fractions $f^{T}$. Since thermal buoyancy is not expected to contribute more than 25% of the total buoyancy power for Earth’s core (Lister and Buffett, 1995), the geodynamo can exist below the threshold for polarity reversals for $Ra_{\ell,H}/Ra^{T}_{\ell}>0.5$ (table 4), which must in turn correspond to large lower-mantle heat flux variations $q^{*}$ of $O(10)$ as per figure 15(d). The self-consistent two-component dynamo model, subject to heterogeneous outer boundary heat flux derived from the seismic shear wave velocity in Earth’s lower mantle, further substantiates this point (see table 5 in Appendix C). The estimated value of $q^{*}$ is regarded as an upper bound because it corresponds to the suppression of the slow MAC waves, at which state the dynamo exhibits reversals consistent with the occasional reversals observed in Earth. Beyond this limit, the dynamo transitions to a multipolar state. For relatively small horizontal buoyancy, the dynamo would be comfortably placed in the dipole-dominated regime without polarity transitions. In short, for nearly invariant vertical Rayleigh numbers of composition and temperature, the magnitude of the horizontal Rayleigh number would determine whether the dynamo operates in a non-reversing dipolar state, a reversing state or a multipolar state.
SL $B^{2}_{peak}$ $f^{T}\%$ $Ra^{C}_{\ell,V}$ $Ra^{T}_{\ell,V}$ $Ra_{\ell,H}$ $\omega_{A,V}^{2}/\omega_{M}^{2}$ $\omega_{A,H}^{2}/\omega_{M}^{2}$ $Ra_{\ell,H}/Ra^{T}_{\ell}$
@@ 4.2.1 A constraint on the lateral heat flux variation in Earth’s lower mantle
Table 4: Calculation of the relative intensity of horizontal buoyancy (last column) in two-component linear magnetoconvection for states where the slow MAC waves disappear. The thermal power ratio $f^{T}$ is defined in (75); $Ra^{C}_{\ell,V}$, $Ra^{T}_{\ell,V}$ are the local vertical compositional and thermal Rayleigh numbers respectively. In addition, $Ra_{\ell,H}$ is the local horizontal Rayleigh number. The resultant local thermal Rayleigh number is $Ra_{\ell}^{T}=Ra^{T}_{\ell,V}+Ra_{\ell,H}$. The parameters are $E_{\eta}=2\times 10^{-5}$ and $t=0.01$.
@@ 5 Concluding remarks
The present study investigates the dipole–multipole transition through the analysis of MHD wave motions in rapidly rotating dynamos subject to inhomogeneous heat flux at the outer boundary. The regime in focus is that of low inertia, where the nonlinear inertial force is small relative to the Coriolis force not only on the length scale of the planetary core but also on the characteristic length scale of convection. The dipole–multipole transition corresponds to the approximate parity between $|\omega_{M}|$ and $|\omega_{A}|$, where $|\omega_{M}|$ is based on the peak field intensity and $|\omega_{A}|$ is a resultant buoyancy frequency derived from the vertical and horizontal (lateral) buoyancy frequencies. The present study focuses on non-axisymmetric outer boundary heat flux, motivated by mantle convection models. That said, as indicated by earlier studies (e.g. Glatzmaier et al., 1999), a high equatorial axisymmetric boundary heat flux can also induce the dipole–multipole transition in rotating dynamos. An axisymmetric positive equatorial heat flux anomaly is analogous to the enhanced equatorial heat flux in a model with homogeneous boundary heat flux (Majumder et al., 2024), where the parity between $|\omega_{M}|$ and $|\omega_{A}|$ in equation (1) causes the disappearance of the slow waves, and in turn, the polarity transition. For the axisymmetric, equatorially anti-symmetric $Y_{1}^{0}$ heterogeneity, the resultant $\beta$ in equation (49) must be evaluated with $\beta_{z}$ taken under a modulus sign in regions where $\partial T/\partial r<0$, such that an increase in $q^{*}$ (or, equivalently, in the magnitude of $\beta_{z}$) leads to an increase in the resultant buoyancy frequency $|\omega_{A}|$. Under this procedure, preferential cooling of either the Northern or the Southern hemisphere has the same outcome, likely inducing polarity transitions. However, the theoretical model in §2 does not predict the total suppression of convection under stable stratification, the analysis of which should be the subject of a future study.
A dipole-dominated, strong-field dynamo with a mean square magnetic field intensity $B^{2}/(2\varOmega\rho\mu\eta)=O(1)$ undergoes a polarity transition through the progressive increase of horizontal buoyancy induced by an equatorially anti-symmetric heat flux heterogeneity at the outer boundary. The transition occurs not by the loss of equatorial symmetry of the convection columns, but by the selective suppression of the slow MAC waves. Polarity reversals lie in a range of horizontal buoyancies between the dipolar and multipolar regimes. A non-axisymmetric, equatorially symmetric heat flux variation does not induce the transition even at high values of the dimensionless heterogeneity $q^{*}$ since the mean temperature gradient at the equator is unaffected for a fixed vertical buoyancy. A boundary heterogeneity consisting of comparable magnitudes of symmetric and anti-symmetric variations induces the polarity transition at a value of $q^{*}$ which is of the same order as that for the transition induced by a purely anti-symmetric variation. Whether the rapid collapse of the axial dipole during reversals may be correlated with the decay of the slow MAC waves when $|\omega_{A}|\approx|\omega_{M}|$ is a problem that requires further investigation.
@@
The present study proposes a dynamical constraint on the upper bound of the heat flux heterogeneity based on the complementarity between the vertical and horizontal buoyancies in suppressing the slow waves, noted in both the linear magnetoconvection and nonlinear dynamo models. In a single-component (thermal) buoyancy-driven dynamo, a dipole solution may be possible for any value in the range 0–0.75 of the relative horizontal buoyancy, $Ra_{\ell,H}/Ra_{\ell}$ (figure 15(c)). However, the fact that compositional buoyancy is dominant in the two-component geodynamo, together with the known order of magnitude of the peak field intensity in the inertia-free limit, suggest $Ra_{\ell,H}/Ra_{\ell}>0.5$, corresponding to $q^{*}$ of $O(10)$ at which an axial dipole can exist. From the point of view of global mantle convection (Olson et al., 2015) and the regional stratification of the outer core (Mound et al., 2019), the maximum variation of heat flux at Earth’s CMB could be higher than the mean superadiabatic heat flux.
For comparable thermal power ratios $f^{T}$, the magnitudes of the upper bound of $q^{*}$ predicted by the two-component magnetoconvection model (§4.1) and obtained from the self-consistent two-component dynamo simulations (Appendix C) are comparable. A lower bound for the heterogeneity may also be obtained from dynamo simulations, based on the minimum variation needed to obtain the hemispherical (east–west) variability in the high-latitude magnetic flux in Earth (Sahoo and Sreenivasan, 2020) and the mantle-induced heat flux heterogeneity at the inner core boundary (Sreenivasan and Gubbins, 2011).
@@
In §4.1, the chosen value of the local compositional Rayleigh number $Ra_{\ell,V}^{C}$ is well supported by observations of the polar core flow (Majumder and Sreenivasan, 2023). That said, the value of $Ra_{\ell,V}^{C}$ in Earth’s core may have varied in the past. Based on the complementarity of buoyancy frequencies in (68), small vertical Rayleigh numbers require large horizontal Rayleigh numbers, and in turn, large $q^{*}$ to achieve reversals. For the present-day Earth, with nearly invariant vertical Rayleigh numbers of composition and temperature, values of the relative horizontal buoyancy much lower than that needed to suppress the slow MAC waves in the core would place the dynamo in a non-reversing dipolar state. This provides a way to explain the existence of geomagnetic superchrons, the long periods in Earth’s history without polarity reversals. That said, even for relative horizontal buoyancies >0.5>0.5, the ratio of symmetric to anti-symmetric heterogeneity in heat flux at Earth’s CMB can vary in geological time. For example, Courtillot and Olson (2007) suggest a relation between superchrons and the formation of mantle plumes, which control the heat flow across the CMB. If a plume forms near the equator, the heat flux at the CMB is dominated by an equatorially symmetric heat flux pattern, promoting a dipole-dominated non-reversing state. On the other hand, if a plume forms far from the equator, the resulting anti-symmetric heat flux pattern suppresses the slow MAC waves and produces a regime conducive to reversals.
@@ Appendix A Dynamo with outer boundary heat flux derived from seismic tomography
Figure 17: Evolution of dipole colatitude with time (measured in units of the magnetic diffusion time) for the dynamo subject to heterogeneous outer boundary heat flux based on the seismic shear wave velocity in Earth’s lower mantle at $q^{*}=13$. The other parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
@@
Figure 18: Section plots at height $z=0.2$ below the equator showing $\omega_{A}^{2}$ (a, c) and $|\omega_{M}^{2}/\omega_{A}^{2}|$ (b, d) for the composite outer boundary heat flux heterogeneity based on the seismic shear wave velocity in Earth’s lowermost mantle. Two values of the heterogeneity are considered, $q^{*}=10$ (a, b) and $q^{*}=13$ (c, d). The other parameters are $Ra_{V}=2500$, $E=1.2\times 10^{-5}$, and $Pm=Pr=1$.
@@ Appendix B Spectral coefficients for the perturbations of temperature and composition
$g_{3}$ $=\frac{\partial^{2}\hat{\Theta}}{\partial{t^{2}}}\big|_{t=0}=a_{2}\bigg(\beta_{y}^{T}\frac{k_{y}}{k_{z}}-\beta_{z}\bigg),$
$g_{4}$ $=\frac{\partial^{3}\hat{\Theta}}{\partial{t^{3}}}\big|_{t=0}=a_{3}\bigg(\beta_{y}^{T}\frac{k_{y}}{k_{z}}-\beta_{z}\bigg),$
$g_{5}$ $=\frac{\partial^{4}\hat{\Theta}}{\partial{t^{4}}}\big|_{t=0}=a_{4}\bigg(\beta_{y}^{T}\frac{k_{y}}{k_{z}}-\beta_{z}\bigg).$
@@
$q_{3}$ $=\frac{\partial^{2}\hat{\gamma}}{\partial{t^{2}}}\big|_{t=0}=a_{2}\beta_{y}^{C}\frac{k_{y}}{k_{z}},$
$q_{4}$ $=\frac{\partial^{3}\hat{\gamma}}{\partial{t^{3}}}\big|_{t=0}=a_{3}\beta_{y}^{C}\frac{k_{y}}{k_{z}},$
$q_{5}$ $=\frac{\partial^{4}\hat{\gamma}}{\partial{t^{4}}}\big|_{t=0}=a_{4}\beta_{y}^{C}\frac{k_{y}}{k_{z}}.$